OpenStudy (anonymous):

In this lesson, you solve linear equations and inequalities. You also explore the difference between identities and contradictions. How do you know when an equation has no real solution, just one solution, two solutions, or an infinite number of solutions? Give as many tips to your classmates as possible. (NOTE: I'm Dyslexic, not getting it.. Poorly worded? Can you at least rephrase it please?)

OpenStudy (anonymous):

And possibly explain the terms better??? My school is not very good, I don't even have a book.

terenzreignz (terenzreignz):

This is pretty broad...

OpenStudy (anonymous):

That is all the information I was given. It's for a discussion board.

terenzreignz (terenzreignz):

How do you know when an equation has one solution.... hmm

terenzreignz (terenzreignz):

Okay, let's play around with a sample equation... \[\Large 2\color{red}x + 3 -7\color{red}x +4 \ = \ 5-3\color{red}x-10 +\color{red}x\] Can you tell from a glance how many solutions this will have?

OpenStudy (anonymous):

There is only one variable, so there can only be one solution, right?

terenzreignz (terenzreignz):

Not necessarily :) How about this one... \[\Large 3\color{red}x +5 - \color{red}x \ = \ -4 +2\color{red}x +9\]

terenzreignz (terenzreignz):

Let's try x = 1 \[\Large 3(1) + 5 -1 = -4 +2(1) + 9\\\Large 7 = 7\] so x = 1 is a solution

terenzreignz (terenzreignz):

Let's try x = 0 \[\Large 3(0) + 5 - 0= -4 +2(0) + 9\\\Large 5 = 5\] x = 0 is also a solution? WTF? LOL

terenzreignz (terenzreignz):

In fact, it doesn't matter what value x takes, it will always be a solution. So this is an equation with infinite solutions :)